# A Nonparametric Bayesian Local-Global Model for Enhanced Adverse Event Signal Detection in Spontaneous Reporting System Data

### v1

@@ 1 Introduction
Adverse events (AEs) related to drugs and other medical products have a profound impact on public health and have received heightened attention over the past decades. While clinical trials remain the gold standard for assessing medical product safety before approval, they have limitations due to factors such as small sample sizes, ethical considerations, restricted timelines, etc., and the critical need for postmarket monitoring is increasingly recognized. For example, severe AEs such as congenital malformations from Thalidomide and liver injury from the Ketoconazole (Lavertu et al.,, 2021) were identified only after these drugs were approved for the U.S. market. To complement clinical trial findings with real-world evidence (RWE), continuous post-marketing surveillance—a cornerstone of contemporary pharmacovigilance—has proven essential. Accordingly, several Spontaneous Reporting System (SRS) databases have been established globally, including VigiBase (World Health Organization), and FAERS and VAERS (U.S. Food and Drug Administration). These databases collect and manage AE data for regulated medical products, such as drugs, medical devices, vaccines, and therapeutic biologics. For exposition, we focus only on drugs hereinafter, though analogous considerations apply to other medical products monitored in SRS databases.
@@
The SRS databases serve as key resources for assessing real-world evidence in postmarket drug safety by cataloging a large number of case safety reports linking drugs to relevant AEs. For example, the FAERS database has compiled over $29$ million reports by 2024. These databases are inherently observational and suffer from selection bias, measurement error, confounding, and missing data, among other factors (Markatou and Ball,, 2014), precluding causal inference. Instead, the focus is on detecting signals: identifying drug-AE pairs that appear disproportionately more often than their null expected counts, derived under the assumption of independence between the drug and the AE using their marginal totals and the dataset’s grand total counts. Association analyses then estimate and assess the significance of signal strengths, quantifying the differences between the observed and (null) expected counts or related quantities defined by parameters of some underlying statistical model.
@@
Numerous SRS data mining and signal detection approaches have been developed over the past decades. Examples include the proportional reporting ratio (PRR) (Evans et al.,, 2001), the reporting odds ratio (ROR) (Rothman et al.,, 2004), formal frequentist methods such as the likelihood ratio test (LRT) and its variants (Huang et al., 2011; Chakraborty et al., 2022; Ding et al., 2020; and the references cited therein), and Bayesian methods such as the Gamma Poisson Shrinker (GPS) (DuMouchel,, 1999), Bayesian confidence propagation neural network (BCPNN) (Bate et al.,, 1998), and Dirichlet process (DP) Poisson Gamma mixture model (Hu et al.,, 2015). Some approaches—e.g., PRR and ROR—employ ad-hoc thresholds on empirical signal strength estimates, while the more formal methods—e.g., LRT and Bayesian approaches—employ rigorous probabilistic tools for statistically guaranteed signal detection, including controlled type I errors and false discovery rates (FDR). Despite these advances, signal detection for rare AEs and newer drugs remains challenging due to their small (marginal total) sample sizes, which leads to small observed and expected report counts in individual drugs, reducing the statistical power to detect the underlying signals. Yet, a large and growing body of research has documented the critical need for reliable pharmacovigilance signal detection precisely under these setups (see, e.g., Poluzzi et al., 2012; Sardella and Belcher, 2018; Raschi et al., 2021, and the references cited therein).
@@
Bayesian methods facilitate principled inference and signal detection for SRS data by modeling drug-AE signal strengths as model parameters with suitable priors and utilizing their posterior distributions. However, prior selection is crucial. An overly restrictive prior—e.g., a single gamma shrinkage prior for all drug-AE signal strength parameters—may cause excessive shrinkage, pulling estimates toward a single common value (Hu et al.,, 2015). Conversely, an overly flexible prior—e.g., independent flat priors for all parameters—may inadequately regularize their estimation, resulting in overfitting and noisy estimates. State-of-the-art Bayesian data mining approaches balance flexibility and regularization using semi/nonparametric shrinkage priors—e.g., a two-component gamma mixture prior (the GPS model of DuMouchel, (1999)) or a DP mixture prior with a base gamma distribution (Hu et al.,, 2015)—for AE signal strengths separately for each drug. This stabilizes signal detection, provided the drugs are treated either all independently or as a single group.
@@
However, real-world SRS databases record drugs with nuanced similarities in their AEs. For instance, various statin drugs, used to lower blood cholesterol and cardiovascular risk, share chemical compositions and function, leading to common AEs like rhabdomyolysis (Montastruc,, 2023). On the other hand, small differences in drug composition and function can cause distinct AEs or variations in AE severity; e.g., myoglobinuria is typically associated with Atorvastatin but not Lovastatin (Seneff,, 2016). Figure 1 shows the pairwise Spearman correlation coefficients for occurrences of 46 common AEs of six statin drugs as cataloged in FAERS (Q3 2014-Q4 2024). These correlations reveal strong yet nuanced between-drug associations: they are consistently above 0.70, ruling out independence, yet many are noticeably below one, suggesting against a single grouping of all drugs.
@@
Figure 1: Spearman correlation matrix for occurrence counts of 46 common AEs (PTs) among statin drugs in the FDA FAERS data (2014 Q1-2020 Q4)
@@
Existing SRS data mining approaches, both Bayesian and frequentist, rarely account for nuanced between-drug associations. Instead, they often assume either independence of drugs (the original LRT, BCPNN, GPS, and DP Poisson-Gamma) or a single group for all drugs (extended LRT of Huang et al., (2011)). These restrictive assumptions limit the statistical power and sensitivity for detecting signals, particularly for rare AEs and newer drugs with small observed and expected counts. Indeed, our extensive simulations (Section 4) show that existing methods suffer from poor sensitivity (after controlling FDR) when detecting signals from datasets containing multiple but weak true signals, rare AEs, and newer drugs.
@@
We address this critical gap with a novel local-global mixture DP-based Bayesian method that balances flexibility and shrinkage in AE signal strengths for each drug while enabling data-adaptive information sharing across drugs based on their associations, enhancing statistical power for signal discovery. In this framework, a local DP models distinct AE signal strength patterns specific to each drug while a global process defines a common pattern across drugs; the final AE signal strengths for a drug are obtained from a probabilistic mixture of these local and global processes. By appropriately controlling the local-global contributions, our approach can flexibly capture complex between-drug associations in SRS data.
@@
Our key contributions are as follows. First, we propose local-global mixture DP models for SRS signal detection (Section 3). Local-global priors are widely used in contemporary statistics (Bhadra et al., 2019 and references cited therein) and have been employed in DP modeling of grouped data, notably via hierarchical DP models (Teh et al.,, 2004). Our framework instead adopts a more transparent two-component mixture of DPs (Antoniak,, 1974) for local-global pattern decomposition. To our knowledge, the use of local-global DP models for principled SRS data mining is novel. We develop efficient MCMC algorithms for model implementation (Section 3.3 and Supplement S.3; software in the Supplement) and leverage FDR-controlled, FNR-optimized hypothesis tests for signal detection (Section 3.4). Second, we provide an intuitive overview of DP-based SRS data mining models to aid broader understanding (Section 3.1). Third, through extensive simulations (Section 4), we show that our methods achieve comparable or superior sensitivity to existing approaches (reviewed in Supplement S.1) when only a few strong signals are present across common AEs but yield substantial sensitivity gains—often twofold or more—when many weak signals exist among rare AEs and newer drugs, all while maintaining strong FDR and type I error control. Finally, our analysis of a real-world statin dataset (Section 5) from the FDA FAERS database demonstrates our methods’ ability to effectively detect AE signals in practice.
@@ 2 SRS DATA, OUR NOTATION, AND PROBABILITY MODELS
An SRS database stores a large number of individual case safety reports that can be summarized as a high-dimensional contingency table cataloging the occurrences of multiple AEs across the rows and various drugs across the columns. Let $I$ and $J$ denote the total number of AEs and drugs under consideration, respectively. Further, let $n_{ij}$ be the total number/count of reported cases for drug-$j$ and AE-$i$; $i=1,\dots,I$ and $j=1,\dots,J$ constituting the contingency table $((n_{ij}))$. Define $n_{i\bullet}=\sum_{j=1}^{J}n_{ij}$, $n_{\bullet j}=\sum_{i=1}^{I}n_{ij}$, and $n_{\bullet\bullet}=\sum_{i=1}^{I}\sum_{j=1}^{J}n_{ij}$.
@@
Signal detection in SRS contingency tables often relies on the observed relative reporting rates, $n_{ij}/E_{ij}$, or a related quantity, as a measure of signal strength. Here, $E_{ij}=n_{i\bullet}n_{\bullet j}/n_{\bullet\bullet}$ represents the null expected count of drug-$j$ and AE-$i$, assuming independence between the drug and AE. Intuitively, if $n_{ij}/E_{ij}\gg 1$, the observed frequency for $(i,j)$ is much higher than the null expected count, suggesting a potential signal. Challenges arise for small $n_{\bullet j}$ (e.g., a new drug) and/or small $n_{i\bullet}$ (rare AE), leading to a small $E_{ij}$ and rendering $n_{ij}/E_{ij}$ to be noisy. Bayesian approaches parameterizing $n_{ij}/E_{ij}$ (or an analogous quantity) using a probability model for $n_{ij}$ can stabilize inference on signal strengths under these setups by incorporating suitable prior distributions that can facilitate principled information sharing across AEs and drugs, leading to improved signal detection.
@@
The Poisson model $n_{ij}\sim\text{Poisson}(\lambda_{ij}E_{ij})$ is commonly used for its analytical simplicity and computational tractability. Here the signal strength parameter $\lambda_{ij}$ parametrizes $n_{ij}/E_{ij}$, with the pair $(i,j)$ considered a signal if $\lambda_{ij}\gg 1$ and non-signal otherwise. When the Poisson assumption is violated, particularly due to an excess of zero counts arising from structural zeros—i.e., drug-AE pairs that are physically or biologically impossible to co-occur—a zero-inflated Poisson (ZIP) model is often preferred. In this framework, $n_{ij}$ is modeled as a mixture of two components: a $\text{Poisson}(\lambda_{ij}E_{ij})$ component for drug-AE pairs that can co-occur and a point mass at 00, denoted by $\delta_{\{0\}}$ (the Kronecker delta), for structural zeros. Importantly, the structural zero positions are typically unknown and inferred from the SRS dataset.
@@
A Bayesian SRS data mining approach begins with such a probability model for $n_{ij}$ (e.g., Poisson or ZIP), yielding a likelihood for the signal strengths $\lambda_{ij}$ that is combined with a prior distribution to derive the posterior distribution of $\lambda_{ij}$ for inference. The marginal distribution of $n_{ij}$, after integrating over the prior, can be more flexible than the likelihood model for $n_{ij}$. Indeed, with an appropriate nonparametric mixture prior, a flexible marginal distribution accommodating ZIP-like behavior can be achieved using only a Poisson likelihood for $n_{ij}$ (see Remark Remark under Section 3). Thus, Bayesian SRS data mining methods are less sensitive to the likelihood choice (Poisson or ZIP) when a flexible prior is used. However, if the zero-inflation probability parameters are of interest, the ZIP model remains useful.
3 NOVEL LOCAL-GLOBAL DP MODELS FOR SRS DATA MINING
We begin in Section 3.1 with a review of the drug-specific DP Poisson model of Hu et al., (2015), outlining its mechanism and providing guidance on the hyperparameters. We then introduce our novel local-global DP framework (Section 3.2) that accommodates between-drug associations and enables effective information sharing across drugs.
3.1 A conceptual overview on the DP Poisson model
Consider the Poisson likelihood $n_{ij}\sim\text{Poisson}(\lambda_{ij}E_{ij})$, with the DP prior (Hu et al.,, 2015):

### v2

@@ 1 Introduction
Adverse events (AEs) related to drugs and other medical products have a profound impact on public health and have received heightened attention over the past decades. While clinical trials remain the gold standard for assessing medical product safety before approval, they have limitations due to factors such as small sample sizes, ethical considerations, restricted timelines, etc., and the critical need for postmarket monitoring is increasingly recognized. For example, severe AEs such as congenital malformations from Thalidomide and liver injury from the Ketoconazole lavertu2021new were identified only after these drugs were approved for the U.S. market. To complement clinical trial findings with real-world evidence (RWE), continuous post-marketing surveillance—a cornerstone of contemporary pharmacovigilance—has proven essential. Accordingly, several Spontaneous Reporting System (SRS) databases have been established globally, including VigiBase (World Health Organization), and FAERS and VAERS (U.S. Food and Drug Administration). These databases collect and manage AE data for regulated medical products, such as drugs, medical devices, vaccines, and therapeutic biologics. For exposition, we focus only on drugs hereinafter, though analogous considerations apply to other medical products monitored in SRS databases.
@@
The SRS databases serve as key resources for assessing real-world evidence in postmarket drug safety by cataloging a large number of case safety reports, each listing the specific drugs taken by a user along with AEs encountered. For example, the FAERS database has compiled over $29$ million reports by 2024. These databases are inherently observational, and the data curated in them suffer from several statistical challenges, including selection bias, measurement error, confounding, missing data, and absence of information on non-events, non-users, and general health characteristics of users, among other factors markatou2014pattern , precluding formal causal inference. In SRS data mining, the focus instead is on detecting signals— identifying drug-AE pairs whose observed counts are disproportionately higher than their null expected counts, the theoretical baseline counts for the pairs under the assumption that the drug and the AE are independent. Then signal detection entails first quantifying the signal strength, typically in terms of some parameterization of the ratio of the observed counts to the null expected counts (or some related quantities) for each drug-AE pair, and then identifying pairs that exhibit sufficiently large signal strengths.
An important consideration in this context is the derivation of these null expected counts for drug-AE pairs. Unlike traditional contingency table-based case-control data analysis, a key challenge here is the absence of information on non-events: SRS databases record only reported drugs and AEs and provide no information on the total number of individuals taking the drugs who did not experience or report a specific AE. This absence means that the true denominators—the actual counts of patients exposed to each drug—are unavailable, precluding direct calculation of absolute risks or incidence rates as in traditional epidemiological designs. Instead, SRS data mining methods rely on statistical comparators (all other reports in the database not involving the drug or AE in question) to approximate marginal total report counts for each drug and AE. These approximated margins are then used to estimate the null expected counts for the drug-AE pairs gravel2023considerations ; kontsioti2023exploring ; zhang2018three ; markatou2014pattern . Despite these limitations, meaningful safety signals can be detected from these databases when appropriate statistical methods are applied, as demonstrated by a growing body of research.
@@
For principled and accurate signal detection from SRS databases, numerous statistical data mining methods have been developed over the past decades. Examples include the proportional reporting ratio (PRR) evans2001use , the reporting odds ratio (ROR) rothman2004reporting , formal frequentist methods such as the likelihood ratio test (LRT) and its variants huang2011likelihood ; chakraborty2022use ; ding2020evaluation , and Bayesian methods such as the Gamma Poisson Shrinker (GPS) dumouchel1999bayesian , Bayesian confidence propagation neural network (BCPNN) bate1998bayesian , and Dirichlet process (DP) Poisson mixture model Hu et al.hu2015signal . Some approaches—e.g., PRR and ROR—employ ad-hoc thresholds on empirical signal strength estimates, while the more formal methods—e.g., LRT and Bayesian approaches—employ rigorous probabilistic tools for statistically guaranteed signal detection, including controlled type I errors and false discovery rates (FDR). Despite these advances, signal detection for rare AEs and newer drugs remains challenging due to their small (marginal total) sample sizes, which leads to small observed and expected report counts in individual drugs, reducing the statistical power to detect the underlying signals. Yet, a large and growing body of researchpoluzzi2012data ; sardella2018pharmacovigilance ; raschi2021value has documented the critical need for reliable early signal detection precisely under these settings involving rare AEs and newer drugs.
@@
In this paper, we argue that an appropriately formulated Bayesian framework can yield important insights for pharmacovigilance under these challenging settings. Bayesian SRS data mining methods enable principled inference and signal detection by modeling drug-AE signal strengths as parameters with suitable priors and utilizing their posterior distributions. Prior specification, however, is crucial for optimal performance. An overly restrictive prior—e.g., a single, common gamma prior for all drug-AE signal strength parameters—may cause excessive shrinkage, pulling estimates toward a single common valuehu2015signal . Conversely, an overly flexible prior—e.g., independent flat priors for all parameters—may inadequately regularize their estimation, resulting in overfitting and noisy estimates. State-of-the-art Bayesian data mining approaches balance flexibility and regularization using semi/nonparametric shrinkage priors, such as a two-component gamma mixture prior (the GPS modeldumouchel1999bayesian ) or a DP mixture prior with a base gamma distributionhu2015signal , for AE signal strengths separately for each drug. These approaches stabilize signal detection; however, they still impose restrictive assumptions, treating drugs in question either entirely independently or as a single homogeneous group.
@@
Real-world SRS databases, however, record drugs with nuanced similarities in their AEs. For instance, various statin drugs, used to lower blood cholesterol and cardiovascular risk, share chemical compositions and function, leading to common AEs like rhabdomyolysis montastruc2023rhabdomyolysis . Yet, due to small differences in drug composition and function they may cause distinct AEs or have variations in AE severity; e.g., myoglobinuria is typically associated with Atorvastatin but not Lovastatin seneff2016statins . Figure 1 shows the pairwise Kendall’s $\tau$ correlation coefficients between six statin drugs with respect to the occurrence counts of 1491 AEs as cataloged in the FAERS database (Q3 2014-Q4 2024; see Section 5 for details on the dataset). These correlations reveal strong yet nuanced between-drug associations: they are consistently above $0.40$ with some as high as $0.68$, ruling out independence, yet many are noticeably around or below $0.50$, suggesting against a single grouping of all drugs.
@@
Figure 1: Kendall’s $\tau$ correlation matrix for occurrence counts of 1491 commonly occurring AEs (PTs) among six statin drugs in the FDA FAERS data (2014 Q1-2020 Q4).
@@
Existing SRS data mining approaches, both Bayesian and frequentist, rarely account for such nuanced between-drug associations. Instead, they often assume either independence of drugs (the original LRT, BCPNN, GPS, and DP Poisson-Gamma) or a single group for all drugs (extended LRT of Huang et al.huang2011likelihood ). These restrictive assumptions limit the statistical power and sensitivity for detecting signals, particularly for rare AEs and newer drugs with small observed and expected counts. Indeed, our extensive simulations (Section 5) show that existing methods suffer from poor sensitivity (after controlling FDR) when detecting signals from datasets containing multiple but weak true signals, rare AEs, and newer drugs.
@@
We address this critical gap with a novel local-global mixture DP-based Bayesian method that balances flexibility and shrinkage in AE signal strengths for each drug while enabling data-adaptive information sharing across drugs with respect to their AE patterns, enhancing statistical power for signal discovery. In this framework, a local DP models distinct AE signal strength patterns specific to each drug while a global process defines a common pattern across drugs; the final AE signal strengths for a drug are obtained from a probabilistic mixture of these local and global processes. By appropriately controlling the local-global contributions, our approach can flexibly capture complex between-drug associations in SRS data.
@@
Our key contributions are as follows. First, we propose local-global mixture DP models for SRS signal detection (Section 4.2). Local-global priors are widely used in contemporary statistics (see Bhadra et al.bhadra2019lasso and references cited therein) and have been employed in DP modeling of grouped data, most prominently through hierarchical DP models teh2004sharing . Such models require handling complex combinatorial partition structures—often represented via Chinese restaurant franchise-style distributions teh2004sharing —making implementation challenging, particularly for sizable datasets. In contrast, our framework adopts a more transparent two-component mixture of DPs antoniak1974mixtures for local-global pattern decomposition. This enables our framework to achieve efficient MCMC implementation using standard distributions while retaining the ability to capture both shared global structure and drug-specific local variations. To our knowledge, the use of local-global DP models for principled SRS data mining is novel. We develop efficient MCMC algorithms for model implementation (Section 4.3 and Supplement S.2; software in the Supplement) and leverage FDR-controlled, FNR-optimized hypothesis tests for signal detection (Section 4.4). Second, we provide an intuitive overview of DP-based SRS data mining models and guidance on prior specification for hyperparameters to aid broader understanding (Section 4.1). Third, through extensive simulations (Section 5), we show that our methods achieve comparable or superior sensitivity to existing approaches (reviewed in Section 3) when only a few strong signals are present across common AEs but yield substantial sensitivity gains—often twofold or more—when many weak signals exist among rare AEs and newer drugs, all while maintaining strong FDR and type I error control. Finally, our analysis of a real-world statin dataset (Section 6) from the FDA FAERS database demonstrates our methods’ ability to effectively detect AE signals in practice from large-scale SRS databases.
@@ 2 SRS DATA, OUR NOTATION, AND PROBABILITY MODELS
An SRS database stores a large number of individual case safety reports, each documenting the list of drugs a user was exposed to and the AEs encountered. These reports can be summarized as a high-dimensional contingency table cataloging the occurrences of all AEs noted in the database across the rows and various drugs across the columns. Further, let $n_{ij}$ be the total number/count of reported cases for drug-$j$ and AE-$i$; $i=1,\dots,I$ and $j=1,\dots,J$ constituting the contingency table $((n_{ij}))$. Define $n_{i\bullet}=\sum_{j=1}^{J}n_{ij}$, $n_{\bullet j}=\sum_{i=1}^{I}n_{ij}$, and $n_{\bullet\bullet}=\sum_{i=1}^{I}\sum_{j=1}^{J}n_{ij}$. As noted in the Introduction, the true total counts of users of a drug or the total number of individuals suffering from a specific AE are not available in SRS data due to the lack of information on non-events and non-users. Instead, SRS data mining methods consider comparator drugs—all drugs other than the drug-$j$ in question—and comparator AEs — all AEs other than the AE $i$ in question—cataloged in the SRS database, and compare the observed prevalence of the drug-$j$, AE-$i$ pair, which is substantially larger than other drug-AE pairs. Using the comparator drugs, and the reported margins $n_{i\bullet}$, $n_{\bullet j}$, and $n_{\bullet\bullet}$ measuring the total number of reported cases specific to AE-$i$, drug-$j$, and all drugs and AEs combined, respectively, they approximate the expected null baseline count $E_{ij}=n_{i\bullet}n_{\bullet j}/n_{\bullet\bullet}$, assuming independence between the drug and the AE. Of note, SRS data mining often focuses on identifying significant AEs of certain groups or subgroups of drugs, constituting only a (small, somewhat homogeneous) subset of all drugs cataloged in the database. Here, an artificial, heterogeneous group of drugs, obtained by collapsing all other drugs in the database, is often created as a ‘baseline’ or ‘reference’ category. This category of “Other drugs” is conceptually too heterogeneous to have any meaningful, unique AEs, but is still included to aid better comparison for disproportionality analysis, and a greater AE (row) total ($n_{i\bullet}$) counts for more accurate $E_{ij}$ estimation. Analogously, a heterogeneous group of ’Other AEs’, obtained by collapsing all AEs outside those of interest (or with very small sample sizes), may be constructed ding2020evaluation .
@@
Signal detection in SRS contingency tables typically relies on the observed relative reporting rates, $n_{ij}/E_{ij}$, or a related quantity, as a measure of signal strength. Intuitively, if $n_{ij}/E_{ij}\gg 1$, the observed frequency for $(i,j)$ is much higher than the null expected count, suggesting a potential signal. Challenges arise for small $n_{\bullet j}$ (e.g., a new drug) and/or small $n_{i\bullet}$ (rare AE), leading to a small $E_{ij}$ and rendering $n_{ij}/E_{ij}$ to be noisy. Bayesian approaches parameterizing $n_{ij}/E_{ij}$ (or an analogous quantity) using a probability model for $n_{ij}$ can stabilize inference on signal strengths under these setups by incorporating suitable prior distributions that can facilitate principled information sharing across AEs and drugs, leading to improved signal detection.
@@
The Poisson model $n_{ij}\sim\text{Poisson}(\lambda_{ij}E_{ij})$ is commonly used for its analytical simplicity and computational tractability. Here the signal strength parameter $\lambda_{ij}$ parametrizes $n_{ij}/E_{ij}$, with the pair $(i,j)$ considered a signal if $\lambda_{ij}\gg 1$ and non-signal otherwise. When the Poisson assumption is violated, particularly due to an excess of zero counts arising from structural zeros—i.e., drug-AE pairs that are physically or biologically impossible to co-occur—a zero-inflated Poisson (ZIP) model is often preferred. In this framework, $n_{ij}$ is modeled as a mixture of two components: a $\text{Poisson}(\lambda_{ij}E_{ij})$ component for drug-AE pairs that can co-occur and a point mass at 0, denoted by $\delta_{\{0\}}$ (the Kronecker delta), for structural zeros. Importantly, the structural zero positions are typically unknown and inferred from the SRS dataset.
@@
A Bayesian SRS data mining approach begins with such a probability model for $n_{ij}$ (e.g., Poisson or ZIP), yielding a likelihood for the signal strengths $\lambda_{ij}$ that is combined with a prior distribution to derive the posterior distribution of $\lambda_{ij}$ for inference. The marginal distribution of $n_{ij}$, after integrating over the prior, can be more flexible than the likelihood model for $n_{ij}$. Indeed, with an appropriate nonparametric mixture prior, a flexible marginal distribution accommodating ZIP-like behavior can be achieved using only a Poisson likelihood for $n_{ij}$ (see Section 4.1). Thus, Bayesian SRS data mining methods are less sensitive to the likelihood choice (Poisson or ZIP) when a flexible prior is used. However, if the zero-inflation probability parameters are of interest, the ZIP model remains useful.
3 A REVIEW OF A FEW EXISTING SRS DATA MINING METHODS
As noted in the Introduction, numerous SRS data mining methods have been proposed in the literature over the past decades. As representatives of the state-of-the-art and competitors of our proposed approach, we focus on four prominent and commonly used SRS data mining methods—both Bayesian and frequentist—that enable (upon appropriate adjustment, as needed) strong false discovery rate controls while permitting reasonable sensitivity for AE signal detection. This section reviews these methods, highlighting the underlying modeling assumptions.
3.1 The Bayesian Propagation Neural Network (BCPNN) method
The BCPNN approach (Bate et al.bate1998bayesian ) assumes that conditional on the grand total count $n_{\bullet\bullet}$, the cell-specific counts $n_{ij}$, the marginal row totals $n_{i\bullet}$, and the marginal column totals $n_{\bullet j}$ each distributed as binomial-beta: (a) $n_{ij}\mid p_{ij}\sim\text{Binomial}(n_{\bullet\bullet},p_{ij}),\ p_{ij}\sim\text{Beta}(\alpha_{ij},\beta_{ij})$, (b) $n_{i\bullet}\mid p_{i\bullet}\sim\text{Binomial}(n_{\bullet\bullet},p_{i\bullet}),\allowbreak\ p_{i\bullet}\sim\text{Beta}(\alpha_{i\bullet},\beta_{i\bullet})$, (c) $n_{\bullet j}\mid p_{\bullet j}\sim\text{Binomial}(n_{\bullet\bullet},p_{\bullet j}),\allowbreak\ p_{\bullet j}\sim\text{Beta}(\alpha_{\bullet j},\beta_{\bullet j})$, where $p_{ij}$ is the cell-specific reporting rate, and $p_{i\bullet}$ and $p_{\bullet j}$ are the marginal row and column reporting rates. We note that the binomial model assumptions for $n_{ij}$, $n_{i\bullet}$ and $n_{\bullet j}$ can be understood as consequences of an independent Poisson model for $n_{ij}$ with certain choices of $p_{ij}$, $p_{i\bullet}$, and $p_{\bullet j}$. Bate et al.bate1998bayesian consider the information content $\text{IC}_{ij}=\log_{2}\left[p_{ij}/(p_{i\bullet}p_{\bullet j})\right]$ as a measure of signal strength for the drug-$j$ AE-$i$ pair, with a higher IC indicating a stronger association. The authors suggest an empirical Bayes estimation of these quantities using $\hat{\alpha}_{ij}=\hat{\alpha}_{i\bullet}=\hat{\beta}_{i\bullet}=\hat{\alpha}_{\bullet j}=\hat{\beta}_{\bullet j}=1$ and $\hat{\beta}_{ij}=1/(E(p_{i\bullet}\mid n_{i\bullet})E(p_{\bullet j}\mid n_{\bullet j}))-1$, with the denominator of $\hat{\beta}_{ij}$ being computed using the posterior distributions $p_{i\bullet}\mid n_{i\bullet}\sim\text{Beta}(\hat{\alpha}_{i\bullet}+n_{i\bullet},\hat{\beta}_{i\bullet}+n_{\bullet\bullet}-n_{i\bullet})$ and $p_{\bullet j}\mid n_{\bullet j}\sim\text{Beta}(\hat{\alpha}_{\bullet j}+n_{\bullet j},\hat{\beta}_{\bullet j}+n_{\bullet\bullet}-n_{\bullet j})$. The consequent (empirical Bayes) posterior distribution of $\text{IC}_{ij}$ can be approximated using Monte Carlo draws or asymptotic normality. A signal is detected at the drug-$j$ and AE-$i$ pair if the 2.5th posterior percentile for $\text{IC}_{ij}$ exceeds 0. While this fixed threshold of 0 is not necessarily guaranteed to control false discoveries, methods to generate data-driven thresholds ensuring a predefined FDR level (e.g., $0.05$) have been proposed.ahmed2010false
3.2 The Gamma Poisson Shrinker (GPS) model
[…]
where $\kappa\in[0,1]$ is the mixing weight hyperparameter, and $\alpha_{1},\beta_{1}>0$, and $\alpha_{2},\beta_{2}>0$ are mixture component-specific gamma hyperparemeters. Here Gamma($\alpha$, $\beta$) denotes a probability distribution with density $f_{\text{gamma}}(y\mid\alpha,\beta)=\frac{1}{\Gamma(\alpha)}\beta^{\alpha}y^{\alpha-1}e^{-\beta y}$, $y>0$; $\alpha,\beta>0$. Intuitively, the two components of the mixture are expected to represent clusters of $\lambda_{ij}$ corresponding to large values (“signals”) and small values (“non-signals”). The posterior distribution of $\lambda_{ij}$ given $n_{ij}$, $E_{ij}$, and the hyperparameters is:
where $\kappa_{ij}=\kappa f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{1},\beta_{1})/[\kappa f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{1},\beta_{1})+(1-\kappa)f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{2},\beta_{2})]$. Under an empirical Bayes approach, the point estimates $\hat{\kappa}$, $\hat{\alpha}_{1}$, $\hat{\beta}_{1}$, $\hat{\alpha}_{2}$, and $\hat{\beta}_{2}$ for the hyperparameters are obtained by maximizing their marginal likelihoods given $n_{ij}$ and $E_{ij}$. The resulting posterior of $\lambda_{ij}$ conditional on these estimated hyperparameters is used for inference and signal detection. Specifically, dumouchel2001empirical define the empirical Bayesian score $E(\log_{2}(\lambda_{ij})\mid n_{ij},E_{ij},\hat{\kappa},\hat{\alpha}_{1},\hat{\beta}_{1},\hat{\alpha}_{2},\hat{\beta}_{2})$ as a point estimate of signal strength for the drug-$j$ and AE-$i$ pair, and also suggest determining the pair to be a signal if the 5th percentile (EB05) of the above conditional posterior for $\lambda_{ij}$ exceeds a specific threshold, commonly 2. Similar to the BCPNN method, the deterministic threshold of $2$ may not necessarily ensure strong FDR control; however, a data-driven threshold controlling FDR to a pre-specified level (e.g., 0.05) may be obtained.ahmed2010false
3.3 DP Poisson method of Hu et al.hu2015signal
Several extensions and generalizations of the two-component mixture prior used in the GPS method have been suggested in the literaturehu2015signal ; zhang2018three ; tan2025flexible . Of these, we focus on one of the most flexible generalizations as proposed in Hu et al.hu2015signal , which employs a Dirichlet process mixture as the prior for $\lambda_{ij}$. Hu et al.hu2015signal consider the same Poisson likelihood $n_{ij}\sim$ Poisson($\lambda_{ij}E_{ij}$) as in GPS, but replace the two-gamma mixture prior for $\lambda_{ij}$ by a DP (infinite) mixture prior: